X Company currently makes a part and is considering buying it from a company that has offered to supply it for $19.18 per unit. This year, per-unit production costs to produce 60,000 units were:

Direct materials $8.10
Direct labor 5.70
Overhead    5.00
Total    $18.80


$240,000 of the total overhead costs were variable; $43,200 of the fixed overhead costs cannot be avoided even if X Company buys the part. If the company buys the part, the resources that are used to make it cannot be used for anything else. Production next year is expected to increase to 64,050 units.

If X Company continues to make the part instead of buying it, it will save
Answer #1

Total overheads of producing 60,000 units = 60,000 x 5

= $300,000

$240,000 of the total overhead costs were variable

Variable overhead per unit = 240,000/60,000

= $4

Hence, fixed overheads = 300,000 - 240,000

= $60,000

Unavoidable fixed overhead = $43,200

Cost of making Cost of buying
Direct material 64,050 x 8.10 = 518,805 0
Direct labor 64,050 x 5.70 = 365,085 0
Variable overheads 64,050 x 4 = 256,200 0
Fixed overheads 60,000 43,200
Cost of supplier 0 64,050 x 19.18 = 1,228,479
Total cost $1,200,090 $1,271,679

If X Company continues to make the part instead of buying it, it will save = Cost of buying - Cost of making

= 1,271,679 - 1,200,090

= $71,589

Kindly comment if you need further assistance. Thanks